OpenStudy (anonymous):

Permutations and combinations. How many Arithmetic Progressions are there in an AP of n terms.

OpenStudy (anonymous):

@KingGeorge Help?

OpenStudy (experimentx):

probably fo consecutive terms \[ \sum_{i=0}^n \binom{n-(3+i)}{1} \]

OpenStudy (experimentx):

with spacing 2 \[ 2 \times \sum_{i=0}^{\left \lfloor n \over 2 \right \rfloor} \binom{\left \lfloor n \over 2 \right \rfloor-(3+i)}{1} \] with spacing 3 \[ 3 \times \sum_{i=0}^{\left \lfloor n \over 3 \right \rfloor} \binom{\left \lfloor n \over 3 \right \rfloor-(3+i)}{1} \]

OpenStudy (experimentx):

looks like i should consider cases ...

OpenStudy (experimentx):

the general case seems difficult .. anyway this serves as bookmark.